OpenStudy (anonymous):

Brian wants to build a ramp that leads to a storeroom 5 meters above the ground. He wants the ramp to be 10 meters long. What should the angle of elevation of the ramp be

OpenStudy (nirmalnema):

 costheta = base / hypo.. 5/10=1/2 theta=60 degree